# Sparse Mean Field Load Balancing in Large Localized Queueing Systems

### 20 Dec 2023 v1

@@
Abstract
Scalable load balancing algorithms are of great interest in cloud networks and data centers, necessitating the use of tractable techniques to compute optimal load balancing policies for good performance. However, most existing scalable techniques, especially asymptotically scaling methods based on mean field theory, have not been able to model large queueing networks with strong locality. Meanwhile, general multi-agent reinforcement learning techniques can be hard to scale and usually lack a theoretical foundation. In this work, we address this challenge by leveraging recent advances in sparse mean field theory to learn a near-optimal load balancing policy in sparsely connected queueing networks in a tractable manner, which may be preferable to global approaches in terms of communication overhead. Importantly, we obtain a general load balancing framework for a large class of sparse bounded-degree topologies. By formulating a novel mean field control problem in the context of graphs with bounded degree, we reduce the otherwise difficult multi-agent problem to a single-agent problem. Theoretically, the approach is justified by approximation guarantees. Empirically, the proposed methodology performs well on several realistic and scalable network topologies. Moreover, we compare it with a number of well-known load balancing heuristics and with existing scalable multi-agent reinforcement learning methods. Overall, we obtain a tractable approach for load balancing in highly localized networks.
keywords:
Load balancing, localized large queueing networks, sparse mean field control, reinforcement learning.
journal: Performance Evaluation
@@ Abstract
The aforementioned decentralized queueing systems with underlying graph structure can be modelled as various variants of multi-agent (partially-observable) Markov decision processes (MDP) Zhang et al. (2021); Chadès et al. (2021). The goal is typically to learn an optimal load balancing policy, which has been done by using numerous available multi-agent reinforcement learning (MARL) approaches, mainly building upon multi-agent MDPs (MA-MDP) Chu et al. (2020); Lin et al. (2021); Liu et al. (2022) and decentralized partially observable MDPs (Dec-POMDP) Guicheng and Yang (2022); Amato and Oliehoek (2015). However, these methods are usually either not highly scalable or are difficult to analyze Zhou et al. (2021). This has resulted in recent popularity of mean field theory for modelling multi-agent systems, via a single representative agent interacting with the mean field (distribution) of all agents Lacker et al. (2023); Cui et al. (2022). Within mean field theory, one differentiates between mean field games (MFG) for agents in a competitive setting Lasry and Lions (2007) and mean field control (MFC) for cooperation Bensoussan et al. (2013). This work focuses on the latter for reducing the overall jobs dropped in the system.
@@
Note that mean field limits have also been used to study load balancing algorithms, such as JSQ and the power-of-$d$ variants (where the load balancer only obtains the state information of $d\ll M$ out of the total available queues) Mitzenmacher (2001), in terms of sojourn time and average queue lengths Winston (1977); Mukherjee et al. (2018); Dawson et al. (2005). The asymptotic analysis of load balancing policies for systems with different underlying random dense graphs has already been studied to show that as long as the degree $d(M)$ scales with the number of servers, the topology does not affect the performance of the Mukherjee et al. (2018), while in our work we consider the sparse case. Graphons are also used to describe the limit of dense graph sequences Lovász (2012) and have been studied with respect to both MFG and MFC Hu et al. (2022); Cui and Koeppl (2022). However, for sparse graphs, the limiting graphon is not meaningful, making it unsuitable for networks whose degree does not scale with system size, see also Caines and Huang (2021); Lovász (2012). MFGs for relatively sparse networks have been studied using Lp-graphs Fabian et al. (2022), but they too cannot be extended to bounded degrees. We believe that bounded-degree graph modelling is necessary to truly represent and analyse large fixed-degree distributed systems to avoid the need for increasingly global interactions and thus difficulties in scaling.
@@
In models where the graph degree is fixed and small, for different topologies such as the ring and torus, the power-of-two policy was studied using pair approximation to analyse its steady state and to show that choosing between the shorter of two local neighbouring queues improves the performance drastically over purely random job allocations Gast (2015). In contrast, in this work, we will also look at a similar model, where the degree $d$ does not scale with the number of agents, and find that one can improve beyond existing power-of-$d$ policies. In order to find scalable load balancing policies, we apply relatively new results from sparse mean field approximations, see Lacker et al. (2023) and references therein. Recently, mean field analysis of the load balancing policies was done using coupling approach but mainly for spatial graphs and only for power-of-$d$ algorithms, see Rutten and Mukherjee (2023). They consider tasks and servers which may be non-exchangeable on random bipartite geometric graphs and random regular bipartite graphs. To the best of our knowledge, MFC on sparse (bounded degree) graphs has neither been analyzed theoretically in its general form, nor has it been applied to queueing systems.
@@
Additionally, algorithms such as JSQ, SED and their power-of-$d$ variant are based on the unrealistic assumption of instantaneous knowledge of the queue states, which is not true in practice, especially in large systems. We remove this assumption by introducing a synchronization (communication) delay $\Delta t$ into our system model. Though in non-local systems it has already been proven that for $\Delta t=0$, JSQ is the optimal load balancing policy, it is also known that as the delay increases, JSQ fails due to the phenomenon of ”herd behaviour” Mitzenmacher (2001), i.e. allocating all packets simultaneously to momentarily empty queues. It has also been shown that as $\Delta t\to\infty$, random allocation is the optimal policy Mitzenmacher (2001). While for the in-between range of synchronization delays in non-sparse queueing systems, a scalable policy has been learned before using mean field approximations Tahir et al. (2022), it only considered fully connected graphs with sampling $d$ neighbours, and not the sometimes more realistic scenario of local, sparsely connected queueing networks. For other works in this direction, we refer to Mukherjee et al. (2018); Zhou et al. (2021); Lipshutz (2019) and references therein.
@@ 1 Queueing System
We consider a system with a set of $\mathcal{N}=\{1,\ldots,N\}$ schedulers/agents and =$\mathcal{M}=\{1,\ldots,M\}$ servers, where each server has its own queue with finite buffer capacity, $B$, see also Figure 1 for an overview. The queues work in a first-in-first-out (FIFO) manner, and servers take the jobs one at a time from their queue, processing them at an exponential rate $\alpha$. Once a job has been processed, it leaves the system forever. Somewhat to the successful power-of-$d$ policies Mitzenmacher (2001), we assume that each scheduler accesses only a limited number (e.g. $d$ out of the $M$) of available queues and can only allocate its arriving jobs to these $d$ accessible queues, with $d\ll M$ and fixed. We assume $M=N$ and associate each server (queue) with one agent, though it is possible to extend the model to varying or even random numbers of queues per agent.
@@
Jobs arrive to the system accordingly to a Markov modulated Poisson process Fischer and Meier-Hellstern (1993) with total rate $\lambda(t)N$, and then are divided uniformly amongst all agents, which is also equivalent to independent Poisson processes at each agent given the shared arrival rate $\lambda(t)$ by Poisson thinning Harremoës et al. (2007). The agent takes the allocation action based on a learned or predetermined, memory-less policy $\zeta$, which considers current queue state information of its $d$ accessible queues. This information is periodically sent by servers to neighbouring agents, such that agents only obtain information on $d$ queues, reducing the amount of messages to be sent. If the job allocation is done to an already full queue, it is dropped and results in a penalty. Similarly, jobs depart from a queue at rate $\alpha$. The goal of the agent is to minimize the overall jobs dropped.
Figure 1: System model for $N=M=4$. Jobs arriving in a certain synchronization time $\Delta t$ are allocated to all agents using Poisson thinning. Each agent can allocate its incoming jobs to its own queue or the other queue it has access to, indicated by the connecting edges between $(s_{1},s_{3})$ and $(s_{2},s_{4})$. These connections will be defined by the chosen graph topology, see Section 2.1.
1.1 Locality and Scalability
@@
Note that in contrast to many other analyzed settings, in this work the $d$ out of $M$ available queues are not sampled randomly for each package, but instead fixed for each agent according to some predefined topology (see also Section 2.1). In other words, we assume a strong concept of locality where agents have access only to a very limited subset of queues, implying also a strong sense of partial observability in the system. The value for $d$ therefore depends on the type of graph topology being considered. Note that an agent always has access to its own queue, as we associate each of the $M$ queues with an agent. Accordingly, our queueing model contains an associated static underlying undirected graph $G=(\mathcal{N},\mathcal{E})$, where $\mathcal{E}\subseteq\mathcal{N}\times\mathcal{N}$ is the set of edges between vertices (agents) $\mathcal{N}\coloneqq\{1,\ldots,N\}$ based on the $d$ queue accessible by the agent. An agent $i$ will have an edge $(i,j)\in\mathcal{E}$ to another agent $j$ whenever they have access to the queues associated to that agent, and vice versa. Therefore, an agent can have a maximum of $d$ edges (neighbours) to other agents. We will denote the set of neighbours of agent $i$ as $N_{i}\coloneqq\{j\in\mathcal{N}\mid(i,j)\in\mathcal{E}\}$.
@@
The motivation to use a graph structure with bounded degree arises from the fact that the corresponding model allows us to find more tractable local solutions for large systems. We need large systems that are regular in some sense, otherwise the system is too heterogeneous to find a tractable and sufficiently regular solution. Therefore, we look at systems where the regularity can be expressed graphically. The regularity condition already includes basic regular graphs such as grids and toruses, but also allows many other random graphs such as the configuration model, see also Section2.1. Here we then apply RL and MFC to find otherwise hard-to-compute near-optimal queueing strategies in this highly local setting. The simplicity of the queueing strategy – instantaneously assigning a packet to one of the neighbouring queues based on periodic and local information – not only allows for fast execution and high scalability, since information does not need to be exchanged for each incoming packet, but also allows for easy addition of more nodes to scale the system to arbitrary sizes.
In the following, we will obtain tractable solutions by first formulating the finite queueing system, then formulating its limiting mean field equivalent as the system grows, and lastly applying (partially-observed) RL and MFC.
1.2 Finite Queueing System
@@
To begin, consider the following system. Each agent $i$ is associated with a local state, $z_{i}\in\mathcal{Z}\coloneqq\{0,1,\ldots,B\}$ and local action $a_{i}\in\mathcal{A}_{i}$. The state $z_{i}$ will be the current queue filling of the $d$ queues from $\mathcal{M}$ accessible to agent $i$, and the set of actions $\mathcal{A}_{i}$ will be the set of these $d$ accessible queues. Hence, we have a finite set of action and state space. The global state of the system is given by $\mathbf{z}=(z_{1},\ldots,z_{N})\in\mathcal{Z}^{N}$. Similarly, the global action is defined as $\mathbf{a}=(a_{1},\ldots,a_{N})\in\mathcal{A}^{N}\coloneqq\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{N}$.
1.2.1 Synchronous system
@@
We want the agents to work in a synchronized manner, e.g. to model communication delays, and also for the servers to send their queue state information to the respective agents once every fixed time interval. To achieve this, we model our system at discrete decision epochs $\{0,\Delta t,2\Delta t,3\Delta t,\ldots\}$, where $\Delta t>0$ is the synchronization delay, the time passing between each decision epoch. The interval $\Delta t$ may be understood as a type of synchronization or update delay, assuming that it takes $\Delta t$ amount of time to obtain updated local information from the servers and update the queueing strategy (e.g. routing table). Note that we may easily adjust $\Delta t$ to approximate continuous time. Using this delay, we can later model our system as a MFC-based RL problem and learn the optimal policy using state-of-the-art RL-based algorithms Sutton and Barto (2018).
1.2.2 Localized policies
@@
1.2.3 Symmetrization
For simplicity of learning, we further assume that any agent may choose to either send to their own queue, or offload uniformly randomly to any of their neighbours, i.e. we consider the actions $\mathcal{A}_{i}=\mathcal{A}=\{0,1\}$ for all agents $i$, of either sending to its own queue (00), or randomly sending to a neighbour ($1$), see Figure 2 for an example visualization. This assumption symmetrizes the model and allows tractable, regular solutions with theoretical guarantees.
@@ 1.2.3 Symmetrization
Indeed, to some extent such symmetrization is necessary to obtain a mean field limit, as otherwise behaviour depends on the ordering of neighbours. Consider a simple one-dimensional line graph where nodes are connected in a straight path. If we use a model that is not symmetric, e.g. if all agents send all their packets to the first of their two neighbours under some ordering of their neighbours, then we obtain different behaviour depending on this ordering, i.e. which neighbour is considered the first and which is the second. Cutting the graph into sets of three, we could define the center node of each set to be the first neighbour for the other nodes, leading to a packet arrival rate of $2\lambda(t)$ at the center node. On the other hand, if we define first neighbours such that there are no overlapping first neighbours between any agents, then any node will have a packet arrival rate of at most $\lambda(t)$. This shows that a certain symmetrization is natural to obtain a scaling limit.
@@
At most, we could consider a solution that anonymizes but still differentiates between neighbors whenever one is fuller than the other, which may be important especially if e.g. queue serving rates are heterogeneous, and can be analogously handled theoretically. However, in our experiments we found that such an assumption complicates training of a RL policy due to the significant addition of action space complexity. Therefore, trading off between policy expressiveness and RL training stability, we use the aforementioned symmetrization.
@@
Figure 2: Visualization of how agents implement their policy. For instance, agent $i=2$ has neighbours $j\in\{1,3\}$. If its action is $a_{2}=0$, it will allocate all its arriving jobs to its own queue $j=2$ (green arrow). In contrast, if $a_{2}=1$ then the arriving jobs are allocated randomly to one of its neighbour $j$ (red arrows).
1.2.4 Dynamical system model
The state of the agent is the state of its associated queue, and is affected by the actions of itself and neighbouring agents. Therefore, at every epoch, $t\in\mathbb{N}$, given the current global state $\mathbf{z}(t)$ and action $\mathbf{a}(t)$, the next local state $z_{i}(t+1)$ of agent $i$ can be calculated independently only using the neighbours’ states $\mathbf{z}_{N_{i}}(t)\equiv\{z_{j}(t)\}_{j\in N_{i}}$ and actions $\mathbf{a}_{N_{i}}(t)\equiv\{a_{j}(t)\}_{j\in N_{i}}$, i.e.
$\mathbb{P}(\mathbf{z}(t+1)\mid\mathbf{z}(t),\mathbf{a}(t))=\prod_{i=1}^{N}\mathbb{P}(z_{i}(t+1)\mid z_{N_{i}}(t),a_{N_{i}}(t))$ (1)

### 22 Mar 2024 v2

@@
Abstract.
Scalable load balancing algorithms are of great interest in cloud networks and data centers, necessitating the use of tractable techniques to compute optimal load balancing policies for good performance. However, most existing scalable techniques, especially asymptotically scaling methods based on mean field theory, have not been able to model large queueing networks with strong locality. Meanwhile, general multi-agent reinforcement learning techniques can be hard to scale and usually lack a theoretical foundation. In this work, we address this challenge by leveraging recent advances in sparse mean field theory to learn a near-optimal load balancing policy in sparsely connected queueing networks in a tractable manner, which may be preferable to global approaches in terms of wireless communication overhead. Importantly, we obtain a general load balancing framework for a large class of sparse bounded-degree wireless topologies. By formulating a novel mean field control problem in the context of graphs with bounded degree, we reduce the otherwise difficult multi-agent problem to a single-agent problem. Theoretically, the approach is justified by approximation guarantees. Empirically, the proposed methodology performs well on several realistic and scalable wireless network topologies as compared to a number of well-known load balancing heuristics and existing scalable multi-agent reinforcement learning methods.
copyright: none
1. Introduction
The increasing demand for computational resources has led to increased research interest in parallel and distributed systems such as data centres and large-scale cloud networks, which can be accurately modelled by large-scale queueing networks due to their stochastic nature (Rodriguez and Guillemin, 2018; Walker et al., 2022). As a result, there is also a renewed interest in studying and designing scalable load balancing algorithms to allow these parallel systems to operate efficiently by reducing queue lengths, minimizing job waiting times, increasing system throughput, etc. (Mishra et al., 2020). This work also focuses on load balancing policies in large networks with the goal of reducing overall job drops in the system. Here, we will consider sparsely connected queueing systems, rather than assuming that all agents can access to all the queues. Many successful centralized and decentralized algorithms have already been proposed where the load balancer (agent) sends jobs to the available parallel $M$ servers (queues). These include (i) join-the-shortest-queue (JSQ) to reduce expected job delay when servers are homogeneous, have infinite buffer space, and independent, identical (i.i.d.), and exponentially distributed job service times (Winston, 1977), (ii) shortest-expected-delay (SED) to minimize the mean response time of jobs when service times are exponential, but with different service rates (Selen et al., 2016), (iii) join-idle-queue (JIQ) to balance the overall system load while reducing communication overhead between load balancers and servers (Lu et al., 2011), and more, see e.g. (der Boor et al., 2022; Mukherjee et al., 2018). However, the first two aforementioned algorithms are asynchronous and assume instantaneous knowledge of queue states, which is not true in practice, especially in large systems. And for JIQ, queues also need to be maintained at the scheduler end which we do not consider in our system model.
Often, large queueing systems are modelled as a decentralized, multi-agent system with an underlying graph topology. Different types of graphs can be used to represent different types of (wireless) network topologies and information structures, for example a cube-connected cycle was used to represent a versatile network of agents in parallel computing (Habibian and Patooghy, 2017). The vertices in the graph represent agents (load balancers) and the edges represent the (possibly wirelessly connected) neighbourhood of each agent, resulting in local states, actions and information exchange. In this work, all agents with access to the same queues form a neighbourhood, and each agent is assumed to only allocate incoming load to queues within their own neighbourhood.
@@ 1. Introduction
The aforementioned decentralized queueing systems with underlying graph structure can be modelled as various variants of multi-agent (partially-observable) Markov decision processes (MDP) (Zhang et al., 2021). The goal is typically to learn an optimal load balancing policy, which has been done by using numerous available multi-agent reinforcement learning (MARL) approaches, mainly building upon multi-agent MDPs (MA-MDP) (Chu et al., 2020; Lin et al., 2021) and decentralized partially observable MDPs (Dec-POMDP) (Guicheng and Yang, 2022). However, these methods are usually either not highly scalable or are difficult to analyze (Zhou et al., 2021). This has resulted in recent popularity of mean field theory for modelling multi-agent systems, via a single representative agent interacting with the mean field (distribution) of all agents (Lacker et al., 2023; Cui et al., 2022). Within mean field theory, one differentiates between mean field games (MFG) for agents in a competitive setting (Lasry and Lions, 2007) and mean field control (MFC) for cooperation (Bensoussan et al., 2013). This work focuses on the latter for reducing the overall jobs dropped in the system.
@@
Note that mean field limits have also been used to study load balancing algorithms, such as JSQ and the power-of-$d$ variants (where the load balancer only obtains the state information of $d\ll M$ out of the total available queues) (Mitzenmacher, 2001), in terms of sojourn time and average queue lengths (Dawson et al., 2005). The asymptotic analysis of load balancing policies for systems with different underlying random dense graphs has already been studied to show that as long as the degree $d(M)$ scales with the number of servers, the topology does not affect the performance (Mukherjee et al., 2018), while in our work we consider the sparse case. Graphons are also used to describe the limit of dense graph sequences (Lovász, 2012) and have been studied with respect to both MFG and MFC (Hu et al., 2022; Cui and Koeppl, 2022). However, for sparse graphs, the limiting graphon is not meaningful, making it unsuitable for networks whose degree does not scale with system size, see also (Caines and Huang, 2021; Lovász, 2012). MFGs for relatively sparse networks have been studied using Lp-graphs (Fabian et al., 2022), but they too cannot be extended to bounded degrees. We believe that bounded-degree graph modelling is necessary to truly represent and analyse large fixed-degree distributed systems to avoid the need for increasingly global interactions and scaling difficulties.
@@
In models where the graph degree is fixed and small, for different topologies such as the ring and torus, the power-of-two policy was studied using pair approximation to analyse its steady state and to show that choosing between the shorter of two local neighbouring queues improves the performance drastically over purely random job allocations (Gast, 2015). In contrast, in this work, we will also look at a similar model, where the degree $d$ does not scale with the number of agents, and find that one can improve beyond existing power-of-$d$ policies. In order to find scalable load balancing policies, we apply relatively new results from sparse mean field approximations, see (Lacker et al., 2023) and references therein. Recently, mean field analysis of the load balancing policies was done using coupling approach but mainly for spatial graphs and only for power-of-$d$ algorithms, see (Rutten and Mukherjee, 2023). They consider tasks and servers which may be non-exchangeable on random bipartite geometric graphs and random regular bipartite graphs. To the best of our knowledge, MFC on sparse (bounded degree) graphs has neither been analyzed theoretically in its general form, nor has it been applied to queueing systems.
@@
Additionally, algorithms such as JSQ, SED and their power-of-$d$ variant are based on the unrealistic assumption of instantaneous knowledge of the queue states, which is not true in practice, especially in large systems. We remove this assumption by introducing a synchronization (communication) delay $\Delta t$ into our system model. Though in non-local systems it has already been proven that for $\Delta t=0$, JSQ is the optimal load balancing policy, it is also known that as the delay increases, JSQ fails due to the phenomenon of ”herd behaviour” (Mitzenmacher, 2001), i.e. allocating all packets simultaneously to momentarily empty queues. It has also been shown that as $\Delta t\to\infty$, random allocation is the optimal policy (Mitzenmacher, 2001). While for the in-between range of synchronization delays in non-sparse queueing systems, a scalable policy has been learned before using mean field approximations (Tahir et al., 2022), it only considered fully connected graphs with sampling $d$ neighbours, and not the sometimes more realistic scenario of local, sparsely connected queueing networks. For other works in this direction, we refer to (Mukherjee et al., 2018; Zhou et al., 2021; Lipshutz, 2019) and references therein.
@@ 2. Queueing System
We consider a system with a set of $\mathcal{N}=\{1,\ldots,N\}$ schedulers/agents and =$\mathcal{M}=\{1,\ldots,M\}$ servers, where each server has its own queue with finite buffer capacity, $B$. The queues work in a first-in-first-out (FIFO) manner, and servers take the jobs one at a time from their queue, processing them at an exponential rate $\alpha$. Once a job has been processed, it leaves the system forever. Somewhat to the successful power-of-$d$ policies (Mitzenmacher, 2001), we assume that each scheduler accesses only a limited number (e.g. $d$ out of the $M$) of available queues and can only allocate its arriving jobs to these $d$ accessible queues, with $d\ll M$ and fixed. We assume $M=N$ and associate each server (queue) with one agent, though it is possible to extend the model to varying or even random numbers of queues per agent. Note that all connections between agents are assumed to be wireless.
@@
Jobs arrive to the system accordingly to a Markov modulated Poisson process (Fischer and Meier-Hellstern, 1993) with total rate $\lambda(t)N$, and then are divided uniformly amongst all agents, which is also equivalent to independent Poisson processes at each agent given the shared arrival rate $\lambda(t)$ by Poisson thinning (Harremoës et al., 2007). The agent takes the allocation action based on a learned or predetermined, memory-less policy $\zeta$, which considers current queue state information of its $d$ accessible queues. This information is periodically sent by servers to neighbouring agents, such that agents only obtain information on $d$ queues, reducing the amount of messages to be sent. If the job allocation is done to an already full queue, it is dropped and results in a penalty. Similarly, jobs depart from a queue at rate $\alpha$. The goal of the agent is to minimize the overall jobs dropped.
2.1. Locality and Scalability
@@
Note that in contrast to many other analyzed settings, in this work the $d$ out of $M$ available queues are not sampled randomly for each package, but instead fixed for each agent according to some predefined topology (see also Section 3.1). In other words, we assume a strong concept of locality where agents have access only to a very limited subset of queues, implying also a strong sense of partial observability in the system. The value for $d$ therefore depends on the type of graph topology being considered. Note that an agent always has access to its own queue. Accordingly, our queueing model contains an associated static underlying undirected graph $G=(\mathcal{N},\mathcal{E})$, where $\mathcal{E}\subseteq\mathcal{N}\times\mathcal{N}$ is the set of wireless edges between vertices (agents) $\mathcal{N}\coloneqq\{1,\ldots,N\}$ based on the $d$ queues accessible by the agent. An agent $i$ will have an edge $(i,j)\in\mathcal{E}$ to another agent $j$ whenever they have access to the queues associated to that agent, and vice versa. Therefore, an agent can have a maximum of $d$ edges (neighbours) to other agents. We will denote the set of neighbours of agent $i$ as $N_{i}\coloneqq\{j\in\mathcal{N}\mid(i,j)\in\mathcal{E}\}$.
@@
The motivation to use a graph structure with bounded degree arises from the fact that the corresponding model allows us to find more tractable local solutions for large systems. We need large systems that are structured in some sense, otherwise the system is too heterogeneous to find a tractable and sufficiently structured solution. Therefore, we look at systems where the structure can be expressed graphically. To avoid confusion, note that the framework is not limited to regular graphs. The framework already includes basic regular graphs such as grids and toruses, but also allows many other irregular random graphs such as the configuration model, see also Section 3.1. Here we then apply RL and MFC to find otherwise hard-to-compute near-optimal queueing strategies in this highly local setting. The simplicity of the queueing strategy – instantaneously assigning a packet to one of the neighbouring queues based on periodic and local information – not only allows for fast execution and high scalability, since information does not need to be exchanged for each incoming packet, but also allows for easy addition of more nodes to scale the system to arbitrary sizes. So in the following, we will obtain tractable solutions by first formulating the finite queueing system, then formulating its limiting mean field equivalent as the system grows, and lastly applying (partially-observed) RL and MFC.
2.2. Finite Queueing System
@@
To begin, consider the following system. Each agent $i$ is associated with a local state, $z_{i}\in\mathcal{Z}\coloneqq\{0,1,\ldots,B\}$ and local action $a_{i}\in\mathcal{A}_{i}$. The state $z_{i}$ will be the current queue filling of the $d$ queues from $\mathcal{M}$ accessible to agent $i$. The set of actions $\mathcal{A}_{i}$ could be the set of these $d$ accessible queues, to which new packets are assigned. Hence, we have a finite set of action and state space. The global state of the system is given by $\mathbf{z}=(z_{1},\ldots,z_{N})\in\mathcal{Z}^{N}$. Similarly, the global action is defined as $\mathbf{a}=(a_{1},\ldots,a_{N})\in\mathcal{A}^{N}\coloneqq\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{N}$.
2.2.1. Synchronous model
@@
We want the agents to work in a synchronized manner, e.g. to model wireless communication delays, and also for the servers to send their queue state information to the respective agents once every fixed time interval. To achieve this, we model our system at discrete decision epochs $\{0,\Delta t,2\Delta t,3\Delta t,\ldots\}$, where $\Delta t>0$ is the synchronization delay, the time passing between each decision epoch. The interval $\Delta t$ may be understood as a type of synchronization or update delay, assuming that it takes $\Delta t$ amount of time to obtain updated local information from the servers and update the queueing strategy (e.g. routing table). Note that we may easily adjust $\Delta t$ to approximate continuous time. Using this delay, we can also model our system as a MFC-based RL problem and learn the optimal policy using state-of-the-art RL-based algorithms (Sutton and Barto, 2018).
2.2.2. Localized policies
@@
2.2.3. Symmetrized packet assignments
@@
For tractable analysis, it is necessary to handle neighbors anonymously, i.e., providing the same probability of assigning to any particular neighbors with the same queue states. Indeed, to some extent such symmetrization is necessary to obtain a useful mean field limit. Otherwise, behaviour depends strongly on the ordering of neighbours, which we cannot model to obtain the mean field limit. Consider a simple one-dimensional line graph where nodes are connected in a straight path. If we allow actions that are not symmetric, e.g. if all agents send all their packets to the first of their two neighbours under some ordering of their neighbours, then we obtain different behaviour depending on this ordering. Cutting the graph into sets of three nodes, we could define the center node of each set to be the first neighbour for the other nodes, leading to a packet arrival rate of $2\lambda(t)$ at the center node. On the other hand, if we define first neighbours such that there are no overlapping first neighbours between any agents, then any node will have a packet arrival rate of at most $\lambda(t)$. Hence, a certain symmetrization is natural for the scaling limit.
@@
At most, we could consider a solution that anonymizes but still differentiates between neighbors whenever one is fuller than the other. This may be important especially if e.g. queue serving rates are heterogeneous, and the according framework is straightforward. However, in our experiments we found that such an assumption complicates training of a RL policy due to the significant addition of action space complexity. Therefore, trading off between policy expressiveness and RL training stability, we assume that any agent may choose to either send to their own queue, or offload uniformly randomly to any of their neighbours. In other words, for simplicity of learning we consider the actions $\mathcal{A}_{i}=\mathcal{A}=\{0,1\}$ for all agents $i$, of either sending to its own queue (00), or randomly sending to a neighbour ($1$), see Figure 1 for an example visualization. This assumption symmetrizes the model and obtains better performance, see Appendix.
@@
Figure 1. Visualization of how agents implement their policy. For instance, agent $i=2$ has neighbours $j\in\{1,3\}$. If its action is $a_{2}=0$, it will allocate all its arriving jobs to its own queue $j=2$ (green arrow). In contrast, if $a_{2}=1$ then the arriving jobs are allocated randomly to one of its neighbour $j$ (red arrows).
2.2.4. Dynamical system model